OpenStudy (anonymous):

Trig/Precal 2 Please help

OpenStudy (anonymous):

\[\frac{ 1+cosu }{ sinu } + \frac{ sinu }{ cosu+1 }\]

OpenStudy (anonymous):

So far I have this: \[\frac{ (1+sinu)^2 + \sin^2u }{ (cosu+1)(sinu) }\] \[\frac{ \sin^2u+2sinu+1+\sin^2u }{ (cosu+1)(sinu) }\]

OpenStudy (anonymous):

I'm supposed to Simplify the trigonometric expression.by the way.

OpenStudy (mathmale):

You are correct in believing that the least common denominator of these two fractions is (sin u)(cos u + 1). But I see problems with your numerators. Would you mind multiplying out (1 + cos u)(1 + cos u)? Please multiply: (sin u)(sin u). Write your results here, please.

OpenStudy (anonymous):

\[ \frac{\cos (u)+1}{\sin (u)}+\frac{\sin (u)}{\cos (u)+1}=\\\frac{\sin ^2(u)+(\cos (u)+1)^2}{\sin (u) (\cos (u)+1)}=\\\frac{\sin ^2(u)+\cos ^2(u)+2 \cos (u)+1}{\sin (u) (\cos (u)+1)}= \] Can you finish it now?

OpenStudy (anonymous):

What is \[ \sin^2(u) + \cos^2(u) \]

OpenStudy (anonymous):

\[ \frac{\sin ^2(u)+\cos ^2(u)+2 \cos (u)+1}{\sin (u) (\cos (u)+1)}=\\\frac{2 \cos (u)+2}{\sin (u) (\cos (u)+1)}=\frac{2}{\sin (u)}=2 \csc (u) \]

OpenStudy (anonymous):

OOH I messed up on the top part, its supposed to be (1+cos^u)^2 Thanks eliassa. You too mathmale. Thanks guys

OpenStudy (anonymous):

YW

OpenStudy (mathmale):

Your showing your work helped me help you. Glad you're making progress!

OpenStudy (anonymous):

Wish I could have gave you both medals haha.